You place a coin with a mass of 0.006 kg 0.0254 meters from the center of the table. There is friction between the coin and the table (μs = 0.6, μk = 0.4). You turn the turntable on and it spins up slowly to 33.3rpm. Will the coin slide off?
Answer #1

Coin will slide of the table, if static friction force on coin is lower than the centripetal force on coin.

Centripetal force on coin will by:

F = m*ac = m*w^2*r

w = 33.3 rpm = 33.3*2*pi/60 = 3.4872 rad/sec

r = 0.0254 m

So,

F = 0.006*0.0254*3.4872^2 = 0.00185 N

Now Friction force on coin will be:

Ff = us*Fn = us*m*g

Ff = 0.6*0.006*9.81

Ff = 0.035 N

Since Ff > F, So coin will not start motion, Coin will not slide off the table.